# Many-Shot In-Context Learning

### v1

@@ 3.4 Reward Modelling: Learning Code Verifiers In-Context
Figure 6: Learning Verifiers In-Context for checking correctness of GSM8K code solutions. Error bars denotes standard error of mean over 3 seeds. See Figure A.22 for a 2-shot prompt example. Best-of-N accuracy. (Left) Average accuracy of top-ranked code solutions (among 4 solutions) based on the verifier score on 200 GSM8K test problems. Best-of-4 selection with 128-shot bridges the gap between Pass@1 accuracy of 77.25% and Pass@4 accuracy of 90% with Gemini 1.0 Pro model. Verifier Confidence. (Right) Conditional Probabilities of the Yes token $\text{I\kern-1.49994ptP}(Yes)$ from the verifier, averaged over all correct and incorrect solutions on test problems.
@@ 4 Many-shot Learning without Human-Written Rationales
Many-shot ICL could potentially be limited by the availability of high-quality human-generated rationales or demonstrations. This is particularly challenging for complex reasoning tasks, such as GPQA (Rein et al., 2023), where human-generated rationales require significant resources and expert knowledge. In this work, we explore two approaches for addressing this issue.
@@
Recent work (Singh et al., 2023; Gulcehre et al., 2023) has introduced Reinforced Self-Training (ReST). In particular, Singh et al. (2023) demonstrated that fine-tuning using model-generated rationales can be more effective than human-generated rationales for problem-solving tasks, and can be viewed as applying expectation-maximization for reinforcement learning. We introduce Reinforced ICL, where we use model-generated rationales for in-context learning. To obtain model-generated rationales, we use a few-shot or zero-shot chain-of-thought prompt as a starting point to sample multiple rationales for each training problem. Then, we select rationales that obtain the correct final answer (we assume access to ground truth final answers), and arrange them into in-context examples containing (problem, rationale) pairs.
@@
Motivated by encouraging results with Reinforced ICL, we go one step further: what if we removed rationales from the many-shot prompt altogether, and prompt the model only with inputs (e.g., problems)? Specifically, the Unsupervised ICL prompt consists of: 1) a preamble, such as, “You will be provided questions similar to the ones below:”, 2) a list of unsolved inputs or problems, and 3) a zero-shot instruction or a few-shot prompt with outputs for the desired output format. See Appendix A.6 for the exact prompts we use.
One hypothesis for how many-shot unsupervised ICL might surpass few-shot learning with human demonstrations is that, when the LLM already possesses the required knowledge to solve a task, any information inserted in the prompt that can narrow down what knowledge is needed for the task becomes helpful. This would be consistent with the view that ICL simply “locates” latent concepts (e.g., math problem-solving) the LLM acquired during pre-training (Xie et al., 2021; Hendel et al., 2023; Wang et al., 2024). As such, any of the prompt components – inputs, outputs, and their mapping – can help locate such concepts. While Unsupervised ICL is broadly applicable, it may not perform well when the outputs are critical for specifying the task, for example, see Figure A.17 for negative results low-resource machine translation.
@@ 4.1 Problem-solving: Hendrycks MATH & GSM8K
We evaluate Reinforced and Unsupervised ICL on the Hendrycks MATH dataset (Hendrycks et al., 2021), which consists of challenging high school competition-level mathematics problems. We use the MATH500 test set from Lightman et al. (2023) to report performance, and our 4-shot MATH prompt for data generation can be found in the Appendix (Figure A.23). For Unsupervised ICL, we append this 4-shot prompt after the unsolved problems (see Figure A.25 for complete prompt). For comparison, we also evaluate ICL with human-written solutions (ground-truth) from the MATH training set.
Our results are shown in the Figure 7 (left). On MATH500, both Reinforced and Unsupervised ICL outperforms ICL with ground-truth solutions in both the few-shot and many-shot regime. For ICL, we observe that the performance improves with more examples in the prompt up to a point, and then declines (with the peak being at about 125 examples). Performance for Reinforced ICL also improves with the number of examples, and reaches a plateau at around 25 examples (while being about 5% higher than ICL), and unlike ICL, we don’t see a significant drop in performance even for a very large number of examples in the context. Notably, many-shot ICL achieves comparable or superior performance when using only problems compared to using problems with solutions. This suggests solutions may be redundant for eliciting problem-solving via in-context learning on this domain, potentially due to the math-related data seen during pretraining.
@@
Singh et al. (2023) found that fine-tuning a model on model-generated solutions from MATH resulted in improved test performance on GSM8K (Cobbe et al., 2021), which has a different distribution of problems than MATH. Here, we investigate whether many-shot ICL also improves transfer performance on GSM8K, indicating an improvement in general problem-solving abilities from in-context learning. Our results in Figure 7 (right) show that this is indeed the case – Reinforced ICL with MATH prompts excels on GSM8K, outperforming ICL with ground truth MATH solutions as well as Unsupervised ICL in the many-shot setting with at least 25 shots. This indicates that model-generated solutions can enable better generalization than just using problems or combining them with ground-truth solutions for ICL.
@@ 4.2 Question Answering: Google-Proof QA (GPQA)
GPQA (Rein et al., 2023) is a multiple-choice QA benchmark, with difficult questions focused on graduate-level reasoning in biology, physics, and chemistry. Following Claude-3 (Anthropic, 2024), we use the diamond split (198 problems) for evaluation. This split focuses on questions where domain experts agree but experts in other domains struggle despite extended effort and internet access. Remaining 250 questions in non-diamond split are used for many-shot ICL with and without human-written rationales. For Reinforced ICL, we use a zero-shot prompt (Figure A.21) to generate multiple rationales on the non-diamond split, solving 129 problems. We also append this zero-shot prompt after the GPQA problems for specifying output format for Unsupervised ICL.
As shown in Figure 8, average test accuracy with ground-truth rationales improves substantially from 5 shots to 125 shots, with the best-performing 125-shot prompt nearly matching the accuracy of the state-of-the-art Claude-3 Opus. However, we do observe a performance degradation with 250 shots. Moreover, Reinforced ICL results indicate that model-generated rationales on GPQA seem to be better than ground-truth rationales up to 25 shots, while resulting in similar performance with more shots. Additionally, Unsupervised ICL does not follow any systematic trend: it sometimes performs better ICL with ground-truth rationales depending on the number of shots, but generally underperforms Reinforced ICL. As noted in Anthropic (2024), GPQA is a small evaluation dataset and has an inherent higher variance across different runs, which might explain the non-systematic trends.
@@ 4.3 Algorithmic and Symbolic Reasoning: Big-Bench Hard
Figure 9: BIG-Bench Hard. Reinforced ICL’s performance over number of shots, averaged across five random seeds. The error bars denote standard deviation. Reinforced ICL outperforms the human-written chain-of-thought prompt for all tasks. Averaged across tasks, chain-of-thought prompting using human-written prompts gets a success rate of 72.1%, while Reinforced ICL gets 83%.
We now evaluate Reinforced ICL on BIG-Bench Hard (Suzgun et al., 2022), a suite of challenging algorithmic reasoning tasks. To reduce the impact of false positives, we select 8 tasks out of 23 in BIG-Bench Hard for which the likelihood of getting a false positive is low: either the answer string is long, or the number of options for each question is large (at least 6). We use the standard 3-shot CoT prompt from Suzgun et al. (2022) to sample 10 rationales per problem from a training set of 150 problem at a temperature of 1.0. We filter the rationales based on final answer correctness and arrange them into prompts containing 3 to 100 (problem, rationale) pairs. We evaluate test performance on a held-out set of 100 problems.
As shown in Figure 9, Reinforced ICL strongly outperforms the standard 3-shot CoT prompt for almost all tasks, and the performance generally improves monotonically with the number of prompts for 7 out of 8 tasks. For a few tasks, Reinforced ICL outperforms the human-written 3-shot prompt even in the 3-shot setting. This result suggests that model-generated rationales can sometimes outperform human-written rationales even when controlling for the amount of data, mirroring the results reported by Singh et al. (2023) for fine-tuning.
5 Analyzing Many-Shot In-Context Learning
In this section, we study how ICL behavior changes from few-shot to the many-shot regime.
@@ 5.1 Overcoming Pre-training Biases with Many-Shot ICL
Figure 10: Overcoming Pre-Training Bias with Many-Shot ICL. (Left) Many-shot ICL overcomes label flips: Test accuracy for sentiment analysis typically improves with more training shots. Flipped and abstract labels eventually approaching the performance of default labels. (Right) Confidence shift in overcoming bias. For flipped and abstract labels, model confidence in its predicted sentiment labels initially drops, then sharply increases with more training shots to similar value, suggesting a period of overcoming pre-training bias. See §5.1 for more details.
While LLMs demonstrate in-context learning of novel tasks, Kossen et al. (2023) suggest that ICL may have difficulty unlearning biases derived from pre-training data. Their experiments, however, focused mainly on few-shot ICL due to LLM context length limitations. Here, we revisit their study using many-shot ICL on the Financial PhraseBank (FP) sentiment analysis dataset (Malo et al., 2014). Like Kossen et al. (2023), we study label relationships that challenge pre-training preferences:
@@
  • Abstract Labels: Semantically-Unrelated labels [‘A’, ‘B’, ‘C’] are used, removing any pre-existing sentiment association (Wei et al., 2023).
@@
For ICL shots, we sample examples from the validation set (with replaced labels) to exhibit the input-label relationship. We report the results in Figure 10 (left). With few shots, test accuracy with replacement labels is much lower than with default labels. This suggests that with few-shot ICL, the model struggles to overcome its pre-existing biases from pre-training. However, as the number of shots increases, performance on flipped and abstract labels dramatically improves, approaching that of default labels.
@@
We also report model confidence on predicted labels in Figure 10 (right). For default labels, confidence in predicted labels steadily increases with more shots. In contrast, for flipped labels, confidence initially drops then sharply increases before reaching a plateau, suggesting a period of overcoming pre-training bias. We posit that the initial drop in performance and confidence in the few-shot regime from 4 to 8 shots may be attributed to the “early ascent” phenomenon (Pan, 2023; Lin and Lee, 2024): a small number of shots may lead to the retrieval of an incorrect skill, which eventually diminishes as task learning takes effect with more shots. Overall, these results indicates that with sufficient in-context examples, LLMs can overcome pre-training biases.
5.2 Learning Non-Natural Language Tasks
@@
We now test many-shot ICL’s ability to learn abstract mathematical functions with numerical inputs. We focus on parity functions and high-dimensional linear classification; these tasks are appealing because their synthetic nature let us test the generality of ICL and determine its applicability beyond the linguistic domain.
@@
Figure 11: In-Context Linear Classification. Test accuracy for 16, 32 and 64 dimensional linear classification problems, averaged across 5 randomly-generated datasets with 25 points per class for each dataset (250 evaluation points total). As we increase the number of shots, the accuracy improves and approximately tracks the performance of the nearest-neighbor baseline trained from scratch on the same data. We use the default implementation of $k$-nearest neighbours (with $k=5$) from scikit-learn (Pedregosa et al., 2011). See Figure A.24 for an example prompt.
Linear Classification in High Dimensions
@@
We first consider linear classification in high dimensions to test the limits of many-shot ICL. Following the setup from Wei et al. (2023), we create datasets with $N$-dimensional inputs vectors and their corresponding labels, where each dimension is a random integer in $[1,1000]$. For each dataset, we randomly sample another $N$-dimensional vector as the decision boundary and a decision threshold (§A.4). We then provide $K$ N-dimensional points above this threshold and $K$ points below that same threshold as in-context exemplars, and the model must determine whether unseen N-dimensional points are above or below the threshold (we do not tell the model the equation or the threshold). While Wei et al. (2023) used only 16 shots per class, we scale ICL up to 2048 shots per class. As shown in Figure 11, many-shot ICL substantially outperforms random-chance accuracy and nearly matches the performance of a strong baseline, namely $k$-nearest neighbors. While $2048$ shots per class perform best when $N=16$, we observe slight accuracy decrease beyond $512$ shots for higher values of $N$ (Figure 11 C, R).
@@ 5.3 Is Many-Shot ICL Sensitive to Example Ordering?
In few-shot in-context learning (ICL), the order of examples within the prompt can significantly impact model performance (Lu et al., 2021; Xiang et al., 2024). Here, we investigate whether such sensitivity to prompt ordering observed in few-shot ICL persists in many-shot scenarios, which remains
@@
largely unexplored. Specifically, we evaluate ten different random orderings of fixed 50 in-context examples from MATH training split and evaluate performance on the held-out MATH500 test set (Lightman et al., 2023).
@@
As Figure 13 reveals, performance varies significantly across different subareas in MATH500. Strikingly, an ordering that that excels in one subarea may perform poorly in another, for example, the best Geometry ordering yields weak results on Number Theory. This fluctuation results in a smaller variation in average performance compared to individual subareas. Overall, these findings highlight a key challenge in ensuring reliable results with many-shot ICL for long-context models.
5.4 NLL May Not Be Predictive of ICL Performance

### v2

@@ 2.4 Reward Modelling with Many-Shot ICL: Learning Code Verifiers
Figure 6: Learning Verifiers In-Context for checking correctness of GSM8K code solutions. Error bars denotes standard error of mean over 3 seeds. See Figure A.5 for a 2-shot prompt. Best-of-N accuracy. (Left) Average accuracy of top-ranked code solutions (among 4 solutions) based on the verifier score on 200 GSM8K test problems. Best-of-4 selection with 128-shot bridges the gap between Pass@1 accuracy of 77.25% and Pass@4 accuracy of 90% with Gemini 1.0 Pro model. Verifier Confidence. (Right) Conditional Probabilities of the Yes token $\text{I\kern-1.49994ptP}(Yes)$ from the verifier, averaged over all correct and incorrect solutions on test problems.
A standard approach to improve LLM reasoning is to use test-time verification [10, 44, 24]. Specifically, an LLM generates multiple candidate solutions for a given problem and a verifier, also known as an outcome reward model, ranks these solutions and selects the best one. Here, we focus on learning such verifiers in-context for code verification.
To create in-context verification examples, we utilize correct and incorrect code solutions in Python generated using Gemini 1.0 Pro [15] on the GSM8K train set. In the prompt, each (problem, solution) pair is appended with the question “Is the solution correct?” followed by the Yes or No token according to ground truth correctness. At inference, we modify each test (problem, solution) pair in the same way and record the logit of the Yes and No tokens (denoted by $L_{Yes}$, $L_{No}$). To compute the verifier score, we use the normalized probability of the Yes token: $\text{I\kern-1.49994ptP}(Yes)=\exp(L_{Yes})/\big(\exp(L_{Yes})+\exp(L_{No})\big)$. We evaluate verifier performance using best-of-4 selection based on the verifier score on 200 problems from GSM8K test set with Gemini 1.0 solutions.
@@ 3 Many-shot Learning without Human-Written Rationales
Many-shot ICL could potentially be limited by the availability of high-quality human-generated rationales or demonstrations. This is particularly challenging for complex reasoning tasks, such as GPQA [52], where human-generated rationales require significant resources and expert knowledge. In this work, we explore two simple approaches for addressing this issue.
@@
Recent work [55] proposed a simplified version of Reinforced Self-Training [19], demonstrating that fine-tuning using model-generated rationales can be more effective than human-generated rationales for problem-solving tasks. Inspired by their work, we introduce Reinforced ICL, where we use model-generated rationales for in-context learning. To do so, we use a zero-shot or few-shot chain-of-thought [62] prompt as a starting point to sample multiple rationales for each training problem. Then, we select rationales that obtain the correct final answer (we assume access to ground truth final answers or correctness checks), and arrange them into in-context examples containing (problem, rationale) pairs.
@@
We now go one step further than Reinforced ICL: what if we removed rationales from the many-shot prompt altogether, and prompt the model only with inputs? Specifically, the Unsupervised ICL prompt consists of: 1) a preamble, such as, “You will be provided questions similar to the ones below:”, 2) a list of unsolved inputs or problems, and 3) a zero-shot instruction or a few-shot prompt with outputs for the desired output format. See §A.2 for the exact prompts we use.
One hypothesis for how many-shot unsupervised ICL might surpass few-shot learning with human demonstrations is that, when the LLM already possesses the required knowledge to solve a task, any information inserted in the prompt that can narrow down what knowledge is needed for the task becomes helpful. This would be consistent with the view that ICL simply “locates” latent concepts (e.g., math problem-solving) the LLM acquired during pre-training [66, 22, 61]. As such, any of the prompt components – inputs, outputs, and their mapping – can help locate such concepts. While Unsupervised ICL is broadly applicable, it may not perform well, for example, when outputs are critical for specifying the task (Figure 9 and A.11).
@@ 3.1 Problem-solving: Hendrycks MATH & GSM8K
We evaluate Reinforced and Unsupervised ICL on Hendrycks MATH [23], which consists of challenging high school competition-level mathematics problems. We use the MATH500 test set from Lightman et al. [33] to report performance, and our 4-shot MATH prompt for data generation can be found in Figure A.6. For Unsupervised ICL, we append this 4-shot prompt after the unsolved problems (see Figure A.8). For comparison, we also evaluate ICL with human-written solutions (ground-truth) from the MATH training set, with the same problems used for many-shot prompts.
Our results are shown in the Figure 7 (left). On MATH500, both Reinforced and Unsupervised ICL outperforms ICL with ground-truth solutions in both the few-shot and many-shot regime. For ICL, we observe that the performance improves with more examples in the prompt up to a point, and then declines (with the peak being at about 125 examples). Performance for Reinforced ICL also improves with the number of examples, and reaches a plateau at around 25 examples (while being about 5% higher than ICL), and unlike ICL, we don’t see a significant drop in performance even for a very large number of examples in the context. Notably, many-shot ICL achieves comparable or superior performance when using only problems compared to using problems with solutions. This suggests solutions may be redundant for eliciting problem-solving via in-context learning on this domain, potentially due to extensive math-related data seen during pretraining.
@@
Singh et al. [55] found that fine-tuning a model on model-generated solutions from MATH resulted in improved test performance on GSM8K [10], which has a different distribution of problems than MATH. Here, we investigate whether many-shot ICL also improves transfer performance on GSM8K, indicating an improvement in general problem-solving abilities from in-context learning. Our results in Figure 7 (right) show that this is indeed the case – Reinforced ICL with MATH prompts excels on GSM8K, outperforming ICL with ground truth MATH solutions as well as Unsupervised ICL in the many-shot setting with at least 25 shots. This indicates that model-generated solutions can enable better generalization than just using problems or combining them with ground-truth solutions for ICL.
@@ 3.2 Question Answering: Google-Proof QA (GPQA)
GPQA [52] is a multiple-choice QA benchmark, with difficult questions focused on graduate-level reasoning in biology, physics, and chemistry. Following Claude-3 [3], we use the diamond split (198 problems) for evaluation. This split focuses on questions where domain experts agree but experts in other domains struggle despite extended effort and internet access. Remaining 250 questions in non-diamond split are used for many-shot ICL with and without human-written rationales. For Reinforced ICL, we use a zero-shot prompt (Figure A.4) to generate multiple rationales on the non-diamond split, solving 129 problems. We also append this zero-shot prompt after the GPQA problems for specifying output format for Unsupervised ICL.
As shown in Figure 8, average test accuracy with ground-truth rationales improves substantially from 5 shots to 125 shots, with the best-performing 125-shot prompt nearly matching the accuracy of the state-of-the-art Claude-3 Opus. However, we do observe a performance degradation with 250 shots. Moreover, Reinforced ICL results indicate that model-generated rationales on GPQA seem to be better than ground-truth rationales up to 25 shots, while resulting in similar performance with more shots. Additionally, Unsupervised ICL does not follow any systematic trend: it sometimes performs better ICL with ground-truth rationales depending on the number of shots, but generally underperforms Reinforced ICL. As noted in Anthropic [3], GPQA is a small evaluation dataset and has an inherent higher variance across different runs, which might explain the non-systematic trends.
@@ 3.3 Algorithmic and Symbolic Reasoning: Big-Bench Hard
Figure 9: BIG-Bench Hard. Reinforced and Unsupervised ICL with varying number of shots, averaged across five random seeds. We evaluate test performance on a held-out set of 100 problems. The error bars denote standard deviation. Reinforced ICL outperforms Unsupervised ICL for all tasks, which in turns outperforms the human-written chain-of-thought (CoT) prompt. Averaged across tasks, CoT prompting using human-written rationales gets a success rate of 72.1%, Unsupervised ICL obtains 77.1%, while Reinforced ICL gets 83%.
We now evaluate Reinforced ICL and Unsupervised ICL on BIG-Bench Hard [56], a suite of challenging algorithmic reasoning tasks. To reduce the impact of false positives, we select 8 tasks out of 23 in BIG-Bench Hard for which the likelihood of getting a false positive is low: either the answer string is long, or the number of options for each question is large (at least 6). For Reinforced ICL, we use the standard 3-shot CoT prompt from Suzgun et al. [56] to sample 10 rationales per problem from a training set of 150 problem at a temperature of 1.0. We filter the rationales based on final answer correctness and arrange them into prompts containing 3 to 100 (problem, rationale) pairs.
As shown in Figure 9, Reinforced ICL strongly outperforms Unsupervised ICL for almost all tasks, which in turn outperforms the standard 3-shot CoT prompt. Performance for Reinforced ICL generally improves monotonically with the number of prompts for 7 out of 8 tasks. These results indicate the Reinforced ICL is a more robust technique than Unsupervised ICL, especially for tasks in which the demonstrations contain crucial information about the task. For a few tasks, Reinforced ICL outperforms the human-written 3-shot prompt even in the 3-shot setting. This result suggests that model-generated rationales can sometimes outperform human-written rationales even when controlling for the amount of data, mirroring the results reported by Singh et al. [55] for fine-tuning.
4 Analyzing Many-Shot ICL
@@ 4.1 Overcoming Pre-training Biases with Many-Shot ICL
Figure 10: Overcoming Pre-Training Bias with Many-Shot ICL. (Left) Many-shot ICL overcomes label flips: Test accuracy for sentiment analysis typically improves with more training shots. Flipped and abstract labels eventually approaching the performance of default labels. (Right) Confidence shift in overcoming bias. For flipped and abstract labels, model confidence in its predicted sentiment labels initially drops, then sharply increases with more training shots to similar value, suggesting a period of overcoming pre-training bias.
While LLMs demonstrate in-context learning of novel tasks, Kossen et al. [30] suggest that ICL may have difficulty unlearning biases derived from pre-training data. Their experiments, however, focused mainly on few-shot ICL due to LLM context length limitations. Here, we revisit their study using many-shot ICL on the Financial PhraseBank (FP) sentiment analysis dataset [40]. Like Kossen et al. [30], we study label relationships that affect pre-training biases:
@@
  • Abstract Labels: We use [‘A’, ‘B’, ‘C’], removing any pre-existing sentiment association [63].
@@
For ICL shots, we sample examples from the validation set (with replaced labels) to exhibit the input-label relationship and report the results in Figure 10. With few shots, test accuracy with replacement labels is much lower than with default labels. This suggests that with few-shot ICL, the model struggles to overcome its pre-existing biases from pre-training. However, as the number of shots increases, performance on flipped and abstract labels dramatically improves, approaching that of default labels. For default labels, confidence in predicted labels steadily increases with more shots, as shown in Figure 10 (right). In contrast, for flipped labels, confidence initially drops then sharply increases before reaching a plateau, suggesting a period of overcoming pre-training bias.
@@
We posit that the initial drop in performance and confidence in the few-shot regime may be attributed to the “early ascent” phenomenon [47, 36]: a small number of shots may lead to the retrieval of an incorrect skill, which eventually diminishes as task learning takes effect in the many-shot regime. Overall, these results indicate that many-shot ICL can overcome pre-training biases.
4.2 Learning Non-Natural Language Tasks: High-Dimensional Functions
@@
We now test many-shot ICL’s ability to learn abstract mathematical functions with numerical inputs, which let us stress test its generality and applicability to possibly unseen tasks.
@@
Figure 11: In-Context Classification. Test accuracy for 16, 32 and 64 dimensional linear classification problems, averaged across 5 randomly-generated datasets with 25 points per class for each dataset (250 evaluation points total). As we increase the number of shots, the accuracy improves and approximately tracks the performance of the nearest-neighbor baseline trained from scratch on the same data. We use the default implementation of $k$-nearest neighbours (with $k=5$) from scikit-learn [48]. See Figure A.7 for an example prompt.
Binary Linear Classification in High Dimensions
@@
Following the setup from Wei et al. [63], we create datasets with $N$-dimensional inputs vectors and their binary class labels, where each dimension is a random integer in $[1,1000]$. See more details in §A.5. While Wei et al. [63] used only 16 shots per class, we scale ICL up to 2048 shots per class. As shown in Figure 11, while $2048$ shots per class perform best when $N=16$, we observe slight accuracy decrease beyond $512$ shots for higher values of $N$ (Figure 11 C, R). Moreover, many-shot ICL substantially outperforms random-chance accuracy and nearly matches the accuracy of a strong baseline, namely $k$-nearest neighbors, indicating that many-shot ICL can implement nearest-neighbour search over inputs. This is reminiscent of induction heads that implement prefix matching over sequences [46], a plausible mechanism for ICL abilities.
@@ 4.3 Is Many-Shot ICL Sensitive to Example Ordering?
In few-shot in-context learning (ICL), the order of examples within the prompt can significantly impact model performance [38, 65]. Here, we investigate whether such sensitivity to prompt ordering observed in few-shot ICL persists in many-shot scenarios, which remains
@@
largely unexplored. Specifically, we evaluate ten different random orderings of fixed 50 in-context examples from MATH training split and evaluate performance on the held-out MATH500 test set [33].
@@
As Figure 13 reveals, performance varies significantly across different subareas in MATH500. Strikingly, an ordering that that excels in one subarea may perform poorly in another, for example, the best Geometry ordering yields weak results on Number Theory. This fluctuation results in a smaller variation in average performance compared to individual subareas. One interesting extension would be to optimize many-shot prompts using frameworks like DSPy [28] that has been successfully applied for optimizing few-shot prompts based a given metric. Overall, these findings highlight a key challenge in ensuring reliable results with many-shot ICL for long-context models.
4.4 Many-Shot ICL vs. Supervised Fine-Tuning
Figure 14: Comparing SFT with Many-Shot ICL on low-resource translation. We plot mean performance across 3 seeds. The standard deviation is between 0.1% to 0.5%. Base model corresponds to 1-shot performance of Gemini 1.5 Pro.
Many-shot ICL could make task-specific fine-tuning less essential or, in some cases, even unnecessary, allowing LLMs to tackle a wider range of tasks without specialization. While supervised fine-tuning (SFT) is the dominant LLM paradigm when making use of hundreds or thousands of examples, it is computationally expensive in terms of training. In contrast, many-shot ICL does not require any training, however it has a larger inference cost, which can be substantially reduced with KV caching [49, 64], which might be available off-the-shelf with context caching [12].
Here, we compare many-shot ICL to full fine-tuning for machine translation (§2.1). We run two sets of experiments: one using 250 examples, and another using the entire dev set (997 examples). Our results in Figure 14 show that SFT and ICL performance is quite close for Bemba, while SFT has a slight edge for Kurdish. Overall, these results demonstrate that many-shot ICL can be a viable alternative to SFT for some tasks.
4.5 Comparing Many-Shot Abilities of Frontier LLMs
Figure 15: Many-shot ICL with GPT-4-Turbo and Claude-3-Opus [3] on low-resource machine translation (§2.1).
The strong many-shot results with Gemini 1.5 Pro raises the question of whether other long-context frontier LLMs also benefit from many-shot ICL. To do so, we evaluate GPT-4-Turbo (128K context length) and Claude-3-Opus [3] (200K context length) on the low-resource translation (§2.1). For both these models, many-shot ICL scales favorably on Bemba but do not exhibit much improvement on Kurdish. Notably, 1.5 Pro starts lower than Claude-3 on Bemba but improves more rapidly, achieving much higher performance at 997 shots. It also outperforms GPT-4 in few-shot learning and improves further with more examples. Overall, these results indicate that frontier LLMs exhibit varying degree of many-shot ICL capability.
4.6 Long-context scaling laws may not predict ICL performance